A bag contains a $100 bill and two $20 bills. A second bag contains 1 gold marble and 2 silver marbles. You get to pick one bill out of the first bag. After this, you get to pick a marble out of the second bag. If you get the gold marble, you get to triple the amount of money you pulled from the first bag. If you get a silver marble, you get to double the amount of money you picked from the first bag. Build a probability model for all different amounts of money that you might win.


1
Expert's answer
2021-10-27T13:11:21-0400

probability of taking $100 from first bag:

"P(\$100)=1/3"

probability of taking $20 from first bag:

"P(\$20)=2/3"


probability of taking gold marble from second bag:

"P(gold)=1\/3"

probability of taking silver marble from second bag:

"P(silver)=2\/3"


probability of getting $40:

"P(\$40)=P(\$20)\cdot P(silver)=2/3\cdot 2/3=4/9"

probability of getting $60:

"P(\$60)=P(\$20)\cdot P(gold)=2/3\cdot 1/3=2/9"

probability of getting $200:

"P(\$200)=P(\$100)\cdot P(silver)=1/3\cdot 2/3=2/9"

probability of getting $300:

"P(\$300)=P(\$100)\cdot P(gold)=1/3\cdot 1/3=1/9"